OpenStudy (juannyebe):

.two forces, one of 2N and the other of magnitude 3N, are applied to the ring of a force table. the direction of both forces are unknown. which best describes the limitations on the magnitiude of resultant ? a)F≤5N b)2N≤F≤3N c)F≥3N d)1N≤F≤5N e)F≤2N

OpenStudy (ljetibo):

I have no idea what a force table is, but the cases you have to consider are: 1) when both forces are acting in the same direction, then the resultant force is F1+F2 2) opposite case of 1. When the two forces are acting in opposite directions then the resultant force is F1-F2. If the question didn't specify "limitation on *magnitude* this case would depend on the order of forces F2-F1 != F1-F2 but magnitude means the absolute value so no worries there. This should give you a clear cut answer.

OpenStudy (jiteshmeghwal9):

because \(F_1 \),\( F_2\) are vectors so their resultant is\[F_1+F_2=\sqrt{F_1^2+F_2^2+2F_1F_2\cos \theta}\] If \(\theta=0^0\) means their resultant is maximum. If \(\theta=180^0\) means their resultant is minimum. Remind u that \(\theta\) is angle b/w forces.

OpenStudy (jiteshmeghwal9):

\(\theta=0^0\) means forces are parallel or acting in same directions. \(\theta=180^0\) means forces are anti-parallel or acting in opposite directions.

OpenStudy (jiteshmeghwal9):

so the value of F varies b/w it s maximum & minimum values.

OpenStudy (irishboy123):

no-one knows generally what a "ring of a force table" is. we can all guess but draw it or summat.